OpenStudy (anonymous):

What is the solution, or what are the solutions, of log9x2 = log925? no solution 5 only –5 only –5 and 5

OpenStudy (danjs):

there are a couple log rules that are always used, like 3 or 4 to be memorized

OpenStudy (danjs):

like the power one, you have log of x^2, power can pull down to front 2 log x

OpenStudy (anonymous):

ok

OpenStudy (danjs):

actually here, both sides are log base 9 of (something) you can remove the logs by 9^[log9(x^2)] = x^2

OpenStudy (anonymous):

would it be the first?

OpenStudy (danjs):

\[\large b^{\log _{b}(x) } = x\] inverse of log function ,

OpenStudy (danjs):

x^2 = 25 then is left

OpenStudy (anonymous):

5

OpenStudy (danjs):

the log functions can only have + values of X the inverse of the log, the exponential function, can have any + or - value for X

OpenStudy (danjs):



OpenStudy (danjs):

overall shape is like that for the logs, the domain is x values greater than zero